In a certain compound %age of an element ( atomic mass 12) is 92.3%. What is its mole fraction?

Grade:11

1 Answers

deepak kumar singh AskiitiansExpert-IITM
13 years ago

dear deeksha sharma,

                                as i think your question is not complete because to calculate mole fraction you have the mixture of two or more than two compounds.....if i consider that we have to find out the mole fraction within the compound then we need empirical formula of that compound or the other element/s present in the compound.

if i consider that this question is complete than lets consider the empirical formula 1:n (only two elements are present).

one element is carbon (at. mass = 12 given in question)

lets choose second element is some X......and the atomic weight is Mx

then molecular weight of compound CXn   = (12+Mxn)

from question:

12 = (12+Mxn)*92.7/100

this gives Mxn = 1 : only one case is possible now that is for hydrogen. and the n=1.

compound would be C2H2 (ethyne)

and mole fraction would be 1/2 =0.5

so u can see it all depend on empirical formula.

so please cross check your question

 

regard

deepak kumar singh